OpenStudy (babynini):

f(1) + 1^2[f(1)]^4 = 18 and f(1) = 2, differentiate

OpenStudy (babynini):

Original f(x) + x^2[f(x)]^4 = 18 and f(1) = 2, find f'(1)

OpenStudy (babynini):

@zepdrix :)

zepdrix (zepdrix):

\[\large\rm f(x) + x^2\cdot f(x)^4 = 18\]

OpenStudy (anonymous):

ooh

zepdrix (zepdrix):

So where you stuck? :U

OpenStudy (anonymous):

now it makes a bit of sense

OpenStudy (babynini):

Everywhere! I think I am doing the chain part incorrectly.

OpenStudy (anonymous):

\[y+x^2y^4=18\] implicit diff

OpenStudy (babynini):

Like.. = f'(x)+4[x^2(f(x))^3 but after that i'm not sure where to even go.

OpenStudy (anonymous):

product rule and chain rule for the second part

zepdrix (zepdrix):

Woops you forgot your product rule! :)

OpenStudy (babynini):

ooh. at the (x^2)[(fx)^4]?

OpenStudy (anonymous):

now i see why \[y+x^3y^4=18\] is much nicer notation

zepdrix (zepdrix):

\[\large\rm f+\color{royalblue}{x^2f^4}=18\]\[\large\rm f'+\color{royalblue}{(x^2)'f^4+x^2(f^4)'}=18\]Ya product rule on the second part :)

zepdrix (zepdrix):

ahhh i keep making typos :c

zepdrix (zepdrix):

\[\large\rm f'+\color{royalblue}{(x^2)'f^4+x^2(f^4)'}=18'\]\[\large\rm f'+\color{royalblue}{(x^2)'f^4+x^2(f^4)'}=0\]

OpenStudy (babynini):

\[f(x)+x^24(f(x))^3+2x(fx)^4=18\]

zepdrix (zepdrix):

Yah maybe a little trouble with chain rule, hmm

zepdrix (zepdrix):

\[\large\rm \frac{d}{dx}f(x)^4\quad=\quad 4f(x)^3\cdot\color{red}{f'(x)}\]

OpenStudy (babynini):

hmm so now what D:

zepdrix (zepdrix):

\[\large\rm f+x^2f^4=18\]Able to get to this point ok? :)\[\large\rm f'+2xf^4+4x^2f^3f'=0\]After that, we'll umm.. oh lemme rewrite it in your notation so we can plug stuff in,\[\large\rm f'(x)+2xf(x)^4+4x^2f(x)^3f'(x)=0\]Now let's evaluate this mess at x=1.

zepdrix (zepdrix):

Becomes something like this, ya?\[\large\rm f'(1)+2f(1)^4+4f(1)^3f'(1)=0\]

zepdrix (zepdrix):

And they told us that \(\large\rm f(1)=2\).

OpenStudy (babynini):

[sorry, pc got stuck. But I am back!]

OpenStudy (babynini):

and now we have f'(1)+32+32f'(1)=0 f'(1)=-32/33

zepdrix (zepdrix):

Mmmm ya I think that's right! :) Imma double check our work real quick.

zepdrix (zepdrix):

yay good job! ୧ʕ•̀ᴥ•́ʔ୨

OpenStudy (babynini):

yayay! and we have a green check mark! thank you so much :)